# Sorting the Babble in Babel: Assessing the Performance of Language Detection Algorithms on the OpenAlex Database

### v2

@@ 3.1 Performance by Language
[…]
@@ 3.3 Simulation Results
As regards to the other subfigures, i.e. those showing scores that take processing speed into account ($\gamma>0$), all give a unanimous picture of procedure performance. First, best-performing procedures of each algorithm follow a single, title-based regime for the whole .$2.5<=\beta<=$ 2 range. But also and perhaps more importantly, CLD2, FastSpell, and especially FastText score markedly higher than all other algorithms, and this outperformance grows substantially as both $\beta$ and $\gamma$ values increase.
@@ 4 Discussion
In addition to the quality of its overall performance, FastText has the advantage of returning as output the probability value of the most probable language. In this sense, a two- or multi-step classification procedure would be quite feasible. For instance, FastText could be run on the titles of all articles indexed in the database, then other procedures that perform better in specific contexts could be run on all cases where the probability/confidence score of the detected language is below a certain threshold. Given the imbalanced nature of the OpenAlex database and the inherent algorithmic tendency to overpredict the majority class, a special effort could then be made to improve prediction for English documents: as mentioned in Céspedes et al., (2025), English documents are overrepresented in the OpenAlex database at a rate of around 9%, which negatively affects the representativity of all other languages in the database to a greater or lesser extent, but also reinforces their underrepresentation in the scholarly and knowledge ecosystem at large. In that perspective and given OpenAlex’s status as the first truly multilingual database, improving the database’s language metadata has social and political implications that go well beyond the functionalities that are normally expected of bibliographic databases.
@@ Appendix A OpenAlex linguistic distribution, grouped by metadata availability
Language Title Title & Abstract Title & Journal Name Title, Abstract & Journal Name
@@ Appendix B Pseudocode description for corpus building procedures
Title T FOREACH article, GET title.
Title & abstract, else title TAeT FOREACH article, GET (title, abstract); IF NO abstract, GET title.
Abstract, else title AeT FOREACH article, GET abstract; IF NO abstract, GET title.
Title & journal name, else title TJeT FOREACH article, GET (title, journal name); IF NO journal name, GET title.
Journal name, else title JeT FOREACH article, GET journal name; IF NO journal name, GET title.
Abstract, else journal name, else title AeJeT FOREACH article, GET abstract; IF NO abstract, GET journal name; IF NO journal name, GET title.
Journal name, else abstract, else title JeAeT FOREACH article, GET journal name; IF NO journal name, GET abstract; IF NO abstract, GET title.
Abstract & journal name, else title AJeT FOREACH article, GET (abstract, journal name); IF (NO abstract OR NO journal name), GET title.
Title & abstract, else title & journal name, else title TAeTJeT FOREACH article, GET (title, abstract); IF NO abstract, GET (title, journal name); IF NO journal name, GET title.
Title & journal name, else title & abstract, else title TJeTAeT FOREACH article, GET (title, journal name); IF NO journal name, GET (title, abstract); IF NO abstract, GET title.
Greedy Greedy FOREACH article, GET(title, abstract, journal name); IF NO journal name, GET (title, abstract); IF NO abstract, GET (title, journal name); IF (NO abstract & NO journal name), GET title.
@@ Appendix C Processing times
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.1.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.1.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.2.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.2.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.3.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.3.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.4.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.4.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.5.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.5.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.6.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.6.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.7.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.7.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.8.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.8.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.9.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.9.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.10.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.10.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.11.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.11.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.12.1 Precision
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect
@@ D.12.2 Recall
Corpus CLD2 CLD3 Detect Language FastSpell FastText LangID Langdetect

### v3

@@ 3.1 Performance by Language
As the different subplots show, procedure performance varies both substantially and differently across the different subplots, with each algorithm or corpus type performing more or less well depending on the specific circumstances of each language subset evaluated. Looking at procedure scores in general, language averages for precision are higher than their recall counterparts for all languages, with the sole exception of articles written in English and Other languages. Comparisons across languages also show that average precision is highest for Russian and Indonesian articles, while the lowest scores belong to articles written in English and Other languages; as for recall, mean procedure scores are highest for French, Italian, and Portuguese articles, and are lowest on average in the case of Korean, Japanese, Chinese, and Russian articles. Overall, the performances obtained by the assessed procedures in all language categories confirm the well-known inverse relationship between precision and recall: use of the two-sided Wetzels Bayes Factor shows that both performance measures are inversely correlated at a rate of $\rho=-.18$, with the observed data supporting the presence of correlation over the null hypothesis at a ratio of 127437 to 1 (Wetzels et al., 2012).
@@ 3.3 Simulation Results
As regards to the other subfigures, i.e. those showing scores that take processing speed into account ($\gamma>0$), all give a unanimous picture of procedure performance. First, best-performing procedures of each algorithm follow a single, title-based regime for the whole $5<=\beta<=2$ range. But also and perhaps more importantly, CLD2, FastSpell, and especially FastText score markedly higher than all other algorithms, and this outperformance grows substantially as both $\beta$ and $\gamma$ values increase.
@@ 4 Discussion
In addition to the quality of its overall performance, FastText has the advantage of returning as output the probability value of the most probable language. In this sense, a two- or multi-step classification procedure would be quite feasible. For instance, FastText could be run on the titles of all articles indexed in the database, then other procedures that perform better in specific contexts could be run on all cases where the probability/confidence score of the detected language is below a certain threshold. Given the imbalanced nature of the OpenAlex database and the inherent algorithmic tendency to overpredict the majority class, a special effort could then be made to improve prediction for English documents: as mentioned in cespedes2025evaluating, English documents are overrepresented in the OpenAlex database at a rate of around 9%, which negatively affects the representativity of all other languages in the database to a greater or lesser extent, but also reinforces their underrepresentation in the scholarly and knowledge ecosystem at large. In that perspective and given OpenAlex’s status as the first truly multilingual database, improving the database’s language metadata has social and political implications that go well beyond the functionalities that are normally expected of bibliographic databases.
@@ Appendix A OpenAlex linguistic distribution, grouped by metadata availability
Language Title
Title &
Abstract
Title &
Journal Name
Title, Abstract &
Journal Name
@@ Appendix B Pseudocode description for corpus building procedures
Title T
FOREACH article, GET title.
Title & abstract,
else title
TAeT
FOREACH article, GET (title, abstract);
IF NO abstract, GET title.
Abstract, else title AeT
FOREACH article, GET abstract;
IF NO abstract, GET title.
Title & journal name,
else title
TJeT
FOREACH article, GET (title, journal name);
IF NO journal name, GET title.
Journal name,
else title
JeT
FOREACH article, GET journal name;
IF NO journal name, GET title.
Abstract,
else journal name,
else title
AeJeT
FOREACH article, GET abstract;
IF NO abstract, GET journal name;
IF NO journal name, GET title.
Journal name,
else abstract,
else title
JeAeT
FOREACH article, GET journal name;
IF NO journal name, GET abstract;
IF NO abstract, GET title.
Abstract & journal name,
else title
AJeT
FOREACH article, GET (abstract, journal name);
IF (NO abstract OR NO journal name), GET title.
Title & abstract,
else title & journal name,
else title
TAeTJeT
FOREACH article, GET (title, abstract);
IF NO abstract, GET (title, journal name);
IF NO journal name, GET title.
Title & journal name,
else title & abstract,
else title
TJeTAeT
FOREACH article, GET (title, journal name);
IF NO journal name, GET (title, abstract);
IF NO abstract, GET title.
Greedy Greedy
FOREACH article, GET(title, abstract, journal name);
IF NO journal name, GET (title, abstract);
IF NO abstract, GET (title, journal name);
IF (NO abstract & NO journal name), GET title.
@@ Appendix C Processing times
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.1.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.1.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.2.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.2.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.3.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.3.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.4.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.4.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.5.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.5.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.6.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.6.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.7.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.7.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.8.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.8.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.9.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.9.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.10.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.10.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.11.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.11.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.12.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@ D.12.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect